OpenStudy (anonymous):

Can someone check my work please? Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal

OpenStudy (anonymous):

Mass of metal = 27.776 g distilled water = 26 mL water Temp int. = 25.2 c metal temp = 100.8 c mixture = 32.6 c

OpenStudy (anonymous):

I got Iron q = mc(delta T) =27.776*4.18*(32.6-25.2) =27.776*4.18*7.4 =859.16

OpenStudy (anonymous):

@dimensionx ?

OpenStudy (anonymous):

@abb0t @terenzreignz @terenzreignz ???????? Please Help..

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@some_someone??

OpenStudy (anonymous):

@Preetha